OpenStudy (anonymous):

algebraic math problem If 5a - 2b = b + 1 = 9, What is the value of a? What i'v got so far is 5a -2b = b + 1 = 9 + 2b + 2b 5a = 3b + 1 = 9 - 1 - 1 5a = 3b = 8

OpenStudy (anonymous):

Am i doing this right?

OpenStudy (owlcoffee):

This is an application of an axiom, which is also called "trasitive" property. The axiom states that if a number "a" is equal to another number "b" but at the same time "b" is equal to a number "c", then "a" is equal to "c": \[a=b=c\] then: \[a=c\] \[b=c\] That's a summarization of it, but hopefully you get the point. so fo the problem: \[5a-2b=b +1=9\] we can view "5a-2b" as "a", "b+1" as "b" and "9" as "c", and it resonates with the structure of the trasitive axiom, so we can apply it: \[5a-2b=9\] \[b+1=9\] So now you have two equations, where one is univarable, so take the second one, solve for "b" and replace it on the first equation.

OpenStudy (anonymous):

Oh wow I did not know about that property, and sadly i have many to learn on this road however thanks so much! my work was: background as you taught me 5a -2b = a b+1 = b 9 = c meaning two equations i have now 5a-2b=9 and b+1=9 and then to solve i took the univariable as suggested and was easier b + 1 = 9 -1 -1 b = 8 therefore i used b to solve for the second equation 5a - 2b = 9 5a - 2(8) = 9 5a - 16 = 9 +16 +16 5a = 25 5a / 5 = 25 / 5 a = 5 Thank you once again :D

OpenStudy (owlcoffee):

Very well done!